kittybasil:

Use the Midpoint Rule with the given value of n to approximate the integral. Round the answer to four decimal places...

Angle:

integral from a to b (b-a)/n = base length a+(base length/2) = x value of first height + base length = x value of next height + base length = etc the first time is (base length / 2) because we want midpoint between a and a+baselength then the points following would simply be base length distance away this is just the steps... (because you didn't give the integral the question is asking about x'D)

Angle:



kittybasil:

O_O WAIT NO WHAT I WASN'T DONE TYPING AAAA

Angle:

lol, I mean, you posted it like half an hour ago and just didn't finish x'D

kittybasil:

Bleh I'm sorry x_x

kittybasil:

OKAY FIRST ONE

kittybasil:

\[\Large{\lim_{n\rightarrow\infty}\int_{0}^{\pi}e^{x}cos{x}dx,~n=4}\]

Angle:

I thought we did this already?

kittybasil:

*checking closed Q records*

kittybasil:

OH DARN IT ALL

Angle:

I clearly remembering having to draw out pi/8 stuff for you

kittybasil:

I'm sorry :'c

Angle:

lol np x'D do you still need help with it?

kittybasil:

well I got as far as this:\[\large{\sum_{i=1}^{n=4}(e^{x_{i}})\cos{x_{i}}\Delta x}\rightarrow(e^{x_{i}})\cos{x_{i}}\left(\frac{\pi-0}{4}\right)\]\[\Large{\rightarrow(e^{x_i})\cos{x_i}\left(\frac{\pi}{4}\right)}\]

Angle:

we're uhhh using the midpoint rule right?

kittybasil:

Yeah

Angle:

ohhhh yeah, you're doing it right nvm x'D

Angle:

the sigma sign should still be in front though, I think

kittybasil:

So all \(\large{x_1}\) are multiples of \(\Large{\frac{\pi}{8}}\) because interval \(\large{\frac{\pi}{4}\div4=\frac{\pi}{8}}\)

kittybasil:

professor said take it off the minute I put in a numerical value idk

Angle:

first x = pi/8 then next point is x = pi/8 + pi/4 = 3pi/8

kittybasil:

yup so I put this as my sum:\[[~f(\pi/8)+f(3\pi/8)+f(5\pi8)+f(7\pi/8)~](\pi/8)\]

Angle:

perfect! :)

kittybasil:

Okay that is good lol

Angle:

wait

kittybasil:

?

Angle:

base length is pi/4

kittybasil:

!! I FORGOT TO CHANGE THAT OFF MY PAPER *really irritated with self* Ok, all good otherwise?

Angle:

ye

kittybasil:

YAY

kittybasil:

Okay so... next one is\[\large{\int_{1}^{3}\frac{2}{1+x^2}dx,~n=5}\]

kittybasil:

Riemann form\[\large{\sum_{i=1}^{5}\frac{2}{1+x_i^2}\Delta x}\rightarrow\frac{2}{1+x_i^2}\left(\frac{3-1}{5}\right)\rightarrow\frac{2}{1+x_i^2}\left(\frac{2}{5}\right)\]

Angle:

yup

kittybasil:

Midpoint formula = \(\large{\frac{\Delta x}{2}=\frac{2/5}{2}=5}\)?

kittybasil:

WAIT NO\[\frac{2/5}{2}=\frac{2}{5}\times\frac{1}{2}=\frac{1}{5}\]yeah?

Angle:

yeah so first x = 1 + 1/5 = 6/5

kittybasil:

ah yeah

Angle:

then second x = 6/5 + 2/5 = 8/5

kittybasil:

so basically...\[x_1=\frac{6}{5}\times\frac{2}{5}i\]

kittybasil:

oops that is supposed to be a plus sign

Angle:

o-o I gues??

Angle:

oh yeah then that's right x'D

kittybasil:

okay that's it for this question... if I have more I'll open up a new one so everyone gets medals fairly 😂